# Wyb. Wyspiańskiego 27, 50-370 Wrocław, Poland

### v1

@@ 2.2 Selected complexity measures
The method was designed to transform the datasets for both classification and regression. Each of those tasks has a predefined set of complexity measures aiming to estimate the difficulty of a given dataset. The initially considered sets of complexity measures are consistent with the sets described by Lorena et al. [14, 15] – resulting in the analysis of 22 metafeatures for classification tasks and 12 metafeatures for regression tasks. Out of the 22 describing the classification, two measures expressing class imbalance (C1 and C2) were dropped since the presented epco does not allow modifying the class proportion. The complete set of measures for each task was then examined to investigate the possibilities of their modifications via feature projections. This resulted in the selection of ten measures for the classification task and four measures for the regression task, summarized in Table 1.
The considered measures were selected in the preliminary single-criteria experiment, where the epco algorithm was aiming to achieve 10 different values of complexity, sampled evenly from a range from 0 to 1. The experiment altered a single synthetic dataset per task. The measures for which a meaningful range of the complexity values was possible to obtain were selected for further many-criteria experiments. The only measures not included despite this observation were C3 and C4 from the regression category. Those two measures showed a strong correlation with C1 and C2 and were significantly more costly to compute.
@@
The target values describing five levels of problem complexity (from easy to complex) were also established based on the preliminary experiment and are presented in Table 2. All selected measures, except C1 and C2 for regression tasks, return higher values for more complex tasks. Hence, the target values for most measures increase when the goal is to make the problem more difficult, but for C1 and C2 – target values decrease. The measure values presented in the table were evenly sampled from a range that was possible to obtain in the preliminary experiment. The presented measure values will serve as targets $\mathcal{T}$ in the experiments.
It is worth mentioning that the possibility of obtaining certain measure values will depend on the source distribution. For example, in the case of a synthetic regression dataset where the output is produced as a linear combination of features, the measures describing the linearity of a dataset (L1 and L2 regression measures) will remain constant despite transforming the data. The measures selected in the preliminary experiment and their values, presented in Table 2, should serve as a starting point for the epco configuration, which can be altered with method hyperparameters to suit a particular type of data.
@@ 3 Experiment design
The presented algorithm was evaluated in a series of computer experiments. The implementation of the epco algorithm in Python, as well as the experimental code, is publicly available as a GitHub repository [^1] https://github.com/JKomorniczak/EPCO. The complexity measures used in the research were calculated using the problexity library [12]. The synthetic data generators and the recognition methods came from the scikit-learn library.
@@ 3.1 Source datasets
For the optimization process, the synthetically generated datasets were used as a source distribution ${X,y}$. To sample the source distribution, a commonly used Madelon data generator (make_classification) for the classification, and a random regression data generator (make_regression) were used. For both of the tasks, the data was described by $20$ features – of which only $2$ were informative for the binary classification, and $10$ for the regression. For classification, the default label noise of $1\%$ was present, whereas for the regression, a Gaussian noise with a standard deviation of $1.0$ was introduced. Both types of datasets consisted of $350$ samples. To stabilize the results, each of the presented experiments was performed in $10$ replications, with source datasets generated based on various random seeds.
@@ 3.2 Algorithm configuration
The configuration of the epco algorithm was common for both evaluated tasks. The optimization was performed in $100$ iterations with $100$ individuals in a population. The crossover ratio was $25\%$, the mutation ratio $10\%$, and they both reduced throughout optimization with a decay factor of $0.7\%$. As mentioned, the number of criteria differed depending on the considered task, resulting in the optimization of $10$ measures for classification and $4$ measures for regression problems.
As a result of optimization, the $|\mathcal{C}|+1$ datasets were constructed and saved, with a $|\mathcal{C}|$ indicating the number of considered criteria. This resulted in $11$ classification datasets and $5$ regression datasets constructed based on each source distribution and for each target complexity level. The generated datasets were a result of transformation using population leaders. Such an operation will result in generating datasets best fitted to each individual of the considered criteria $C_{i}$ and to the sum of all criteria $\Sigma_{C}$.
@@ 3.3 Recognition methods and quality metrics
After generating the datasets with various complexities according to the selected measures, the final experiment intended to directly assess the classification and regression qualities on the generated data using state-of-the-art methods. The evaluated classifiers included: k-Nearest Neighbors (knn), Decision Tree (dt), Gaussian Naive Bayes (gnb), Multilayer Perceptron (mlp), and Support Vector Machine (svm). Since the generated datasets were balanced, the classification experiment used the accuracy score in the evaluation.
The evaluated regression methods included: k-Nearest Neighbors Regressor (knr), Decision Tree Regressor (dtr), Bayesian Ridge Regressor (brr), Multilayer Perceptron Regressor (mlpr), and Support Vector Regressor (svr). In the regression experiments, the measure of Mean Absolute Error (mae) was used as a quality metric.
The implementation of all classifiers and regression methods came from scikit-learn and used the default parametrization. Both experiments used 5-times repeated 2-fold cross-validation.
@@ 4 Experiment results
This section will present the results of experiments, both evaluating the epco optimization abilities and the recognition results for classification and regression datasets. The datasets were generated using feature projections established in the complexity measures’ optimization towards five target complexities.
@@ 4.1 Classification
Table 3 presents the complexities of the result datasets achieved in the optimization using epco. The results show average complexities for the projection with the best fitness in the sum of the criteria ($\Sigma_{C}$) – which is the default result of the algorithm. The value in parentheses can be interpreted as an error – the distance from the target to the obtained value.
The results show that most of the target metric values were possible to obtain. The largest error can be seen for T1 measure across all target difficulties and for F1, F3, and ClsCoef measures for extremes of possible target ranges (easy and complex tasks). The lack of possibility to obtain measures on a lower or upper end of the selected range may be the result of considering many criteria in the optimization. The values that were possible to achieve for a single metric in the preliminary experiment – based on which the ranges were selected – may become unreachable when a certain value of another measure is considered equally important. Even though some differences from the targets are observed, the direct values of complexities increase with the expected difficulty for most of the measures.
Another experiment for classification used a larger number of datasets returned by the epco method (the total of $550$ constructed based on $11$ population leaders) and evaluated the recognition quality using five classification methods. The results of classification accuracy obtained in this experiment are presented in Figure 4. The columns present the achieved scores for specific classifiers, with box colors related to their average quality. Reds correspond to the higher accuracy and blues to the lower. The bars represent the results for various requested difficulties from easy to complex.
Figure 4: The results of the classification accuracy achieved by the evaluated classifiers for datasets transformed with epco towards various difficulty levels. Larger complexity of the data resulted in lower accuracy scores for all evaluated classifiers.
The results of this experiment indicate that the datasets with higher complexity are more difficult in terms of the recognition task. This is observed for all of the evaluated classifiers. Interestingly, the classification quality seems to linearly degenerate when increasing the complexity. Since the target values of the complexity measures used in the epco algorithm were also sampled using a linear function in the selected ranges, one may expect that the classification quality is proportional to the complexity measure value.
@@ 4.2 Regression
Table 4 presents the average complexities of the datasets generated using the epco algorithm for the regression task. For regression problems, only 4 measures were used. Since the C1 and C2 measures for regression tasks return lower values for more complex problems, their values decrease when the expected difficulty of the dataset is intended to rise. Similarly to the table presenting results for classification, measures were calculated for the individual best fitted to the sum of the criteria ($\Sigma_{C}$).
@@
In the case of regression tasks, some measures show a stable decrease (C1 and C2) or increase (S1 and S2) when raising the target difficulty. C1 measure was the easiest to optimize – its fitness in parentheses describes the lowest differences from the target. C2 for the easy target is the only case when the optimization error exceeded 0.1. However, as mentioned, targets lying on the extremes of possible ranges may be more difficult to obtain.
@@
Figure 5: The results of the mean absolute error obtained by the evaluated regression methods for datasets transformed with epco towards target difficulties. More challenging problems correspond to higher recognition errors.
In the case of the regression task, one can notice the increase in errors obtained by the evaluated methods when the difficulty of the data increases. The errors differ significantly according to the target complexity in the case of the knr and dtr methods. In the case of mlpr and svmr, the differences in quality are also visible – highlighted by the changes in colors of the boxes – however, the range of obtained mae is smaller compared to other methods. In the case of brr, one can mostly notice the increasing errors of outliers indicated by individual points. The low errors of this method may be a result of the generation of the original data using the make_regression function, describing a linear relation between features and a target.
@@ 5 Conclusions
As a tool to alter the complexity of synthetically generated datasets, a set of 10 complexity measures for classification and 4 complexity measures for regression tasks were optimized in the proposed Evolutionary Projection-based Complexity Optimization approach. The proposed generic algorithm searches the space of feature projection matrices, which are used to transform the source data distribution towards a target complexity profile. The method was used to generate $550$ classification datasets and $250$ regression datasets with five different target difficulty levels. Experiments evaluating state-of-the-art classifiers and regression methods showed a decrease in classification accuracy and an increase in regression errors when faced with datasets of higher complexity.
The presented algorithm used a custom genetic approach to many-criterion optimization, where individuals in the population were sorted according to fitness in the considered criteria and their sum. This simple approach offered promising optimization results, especially considering a small population and a significant number of up to ten considered criteria. Future works may utilize more sophisticated optimization strategies, where the final individuals of the population would form a more continuous pareto frontier of the optimized complexity measures.

### v2

@@ 2.2 Selected complexity measures
The method was designed to transform the datasets for both classification and regression. Each of those tasks has a predefined set of complexity measures estimating the difficulty of a given dataset. The initially considered sets of complexity measures are consistent with those described by Lorena et al. [15, 16], resulting in the analyses of 22 metafeatures for classification tasks and 12 for regression tasks. Out of the 22 designed for classification, two measures expressing class imbalance (C1 and C2) were dropped since the presented epco does not allow modification of the class proportions. The complete set of measures for each task was then examined to investigate the possibility of modifying them via feature projections. This resulted in the selection of ten measures for the classification task and four measures for the regression task, summarized in Table 1.
The considered measures were selected in the preliminary single-criteria experiment, where the epco algorithm aimed to achieve different values of complexity from a range $[0:1]$. The experiment altered a single synthetic dataset per task. The measures for which a meaningful range of complexity values could be obtained were selected for further many-criteria experiments. The only measures not included, despite this observation, were C3 and C4 from the regression category. These two measures showed a strong correlation with C1 and C2 and were significantly more costly to compute.
@@
The target values describing five levels of problem complexity (from easy to complex) were also established based on the preliminary experiment and are presented in Table 2. All selected measures, except C1 and C2 for regression tasks, return higher values for more complex tasks. Hence, the target values for most measures increase when the goal is to make the problem more difficult, but for C1 and C2, the target values decrease. The measure values presented in the table were evenly sampled from a range established in the preliminary experiment. The presented measure values will serve as targets $\mathcal{T}$ in the experiments.
It is worth noting that the possibility of obtaining specific measure values depends on the source distribution. For example, in a synthetic regression dataset where the label is a linear combination of features, the measures of linearity (L1 and L2) will remain constant regardless of the data transformation. The measures selected in the preliminary experiment and their values, presented in Table 2, should serve as a starting point for the epco configuration, which can be altered with method hyperparameters to suit a particular type of data.
@@ 3 Experiment design
The presented algorithm was implemented in Python and evaluated in a series of computer experiments. The code is publicly available [^1] https://github.com/JKomorniczak/EPCO. The complexity measures used in the research were calculated using the problexity library [13]. The implementation of synthetic data generation methods and recognition methods came from scikit-learn and used the default hyperparameters. All experiments used 5-times repeated 2-fold cross-validation.
@@ 3.1 Source datasets
For the optimization process, the synthetically generated datasets were used as a source distribution $\{X,y\}$. For that purpose, we used the Madelon data generator (make_classification) and the random regression data generator (make_regression). For both tasks, the data were described by $20$ features, of which $2$ were informative for the binary classification and $10$ for the regression. For classification, the default label noise of $1\%$ was present, whereas for the regression, Gaussian noise with a standard deviation of $1.0$ was introduced. Both types of datasets consisted of $350$ samples. To stabilize the results, each of the presented experiments was performed in $10$ replications, with source datasets generated based on various random seeds.
@@ 3.2 Algorithm configuration
The configuration of the epco algorithm was common for both evaluated tasks. The optimization was performed over $100$ iterations with $100$ individuals in the population. The crossover ratio was $25\%$, the mutation ratio $10\%$, and both were reduced throughout optimization with a decay factor of $0.7\%$. As mentioned, the number of criteria varied depending on the task, resulting in the optimization of $10$ measures for classification and $4$ for regression.
As a result of the optimization, $|\mathcal{C}|+1$ datasets were constructed and saved, where $|\mathcal{C}|$ denotes the number of considered criteria. This resulted in $11$ classification datasets and $5$ regression datasets constructed based on each source distribution and for each target complexity level. The generated datasets were a result of transformation using population leaders. Such an operation will result in generating datasets best fitted to each individual of the considered criteria $C_{i}$ and to the sum of all criteria $\Sigma_{C}$.
@@ 3.3 Recognition methods and quality metrics
After generating datasets with various complexities according to the selected measures, the final experiment aimed to directly assess the classification and regression qualities of the generated data using state-of-the-art methods. The evaluated classifiers included k-Nearest Neighbors (knn), Decision Tree (dt), Gaussian Naive Bayes (gnb), Multilayer Perceptron (mlp), and Support Vector Machine (svm). Since the generated datasets were balanced, the classification experiment used accuracy as the evaluation metric.
The evaluated regression methods included k-Nearest Neighbors Regressor (knr), Decision Tree Regressor (dtr), Bayesian Ridge Regressor (brr), Multilayer Perceptron Regressor (mlpr), and Support Vector Regressor (svr). In the regression experiments, the Mean Absolute Error (mae) was used as a quality metric.
@@ 4 Experiment results
This section presents the results of experiments evaluating both the epco optimization capabilities and the recognition performance on classification and regression datasets. The datasets were generated using feature projections derived from the optimization of complexity measures towards five target complexities.
@@ 4.1 Classification
Table 3 presents the complexities of the output datasets after optimization using epco. The results show average complexities for the projection with the best fitness in the sum of the criteria ($\Sigma_{C}$) – which is the default result of the algorithm. The value in parentheses can be interpreted as an error – the distance from the target to the obtained value.
The results show that most of the target metric values were possible to obtain. The largest errors are seen for the T1 measure across all target difficulties, and for the F1, F3, and ClsCoef measures at the extremes of the possible target ranges (easy and complex tasks). The inability to obtain measures at the lower or upper ends of the selected range may result from the optimization considering multiple criteria. The values that were achievable for a single metric in the preliminary experiment – based on which the ranges were selected – may become unreachable when another measure is considered equally important. Even though some differences from the targets are observed, the direct values of complexities increase with the expected difficulty for most measures.
Figure 4: The results of the classification accuracy for datasets transformed with epco towards various difficulty levels. The larger complexity of the data resulted in lower accuracy scores for all evaluated classifiers.
Another experiment for classification used a larger number of datasets returned by the epco method (a total of $550$ constructed based on $11$ population leaders) and evaluated the recognition quality using five classification methods. The results of classification accuracy obtained in this experiment are presented in Figure 4. The columns show the achieved scores for specific classifiers, with box colors indicating their average quality. Reds correspond to the higher accuracy, and blues correspond to the lower scores. The bars represent the results for various requested difficulties – from easy to complex.
The results of this experiment indicate that the datasets with higher complexity are more difficult in terms of the recognition task. This is observed for all the evaluated classifiers. Interestingly, classification quality seems to degrade linearly as complexity increases. Since the target values of the complexity measures used in the epco algorithm were also sampled using a linear function in the selected ranges, one may expect that the classification quality is proportional to the value of the complexity measure.
@@ 4.2 Regression
Table 4 presents the average dataset complexity generated by the epco algorithm for the regression task. For regression problems, only 4 measures were used. Since the C1 and C2 measures for regression tasks return lower values for more complex problems, their values decrease when the expected difficulty of the dataset is intended to increase. Similarly to the classification results table, measures were calculated for the individual that best fit the sum of the criteria ($\Sigma_{C}$).
@@
In the case of regression tasks, some measures show a stable decrease (C1 and C2) or increase (S1 and S2) when raising the target difficulty. The C1 measure was the easiest to optimize – its fitness in parentheses describes the smallest differences from the target. C2 for the easy target is the only case in which the optimization error exceeded 0.1. However, as mentioned, targets lying at the extremes of the possible range may be more challenging to obtain.
Figure 5: The results of the mean absolute error obtained by the evaluated regression methods for datasets transformed with epco towards target difficulties. More challenging problems are associated with higher recognition errors.
@@
In the regression task, one can observe an increase in errors for the evaluated methods as the data difficulty increases. The errors differ significantly depending on the target complexity in the case of the knr and dtr methods. In the case of mlpr and svmr, the differences in quality are also visible – highlighted by the changes in the colors of the boxes – however, the range of obtained mae is smaller compared to the remaining methods. In the case of brr, one can mostly notice the increasing errors of outliers indicated by individual points. The low errors of this method may result from the generation of the original data using the make_regression function, which models a linear relationship between features and a target.
@@ 5 Conclusions
To alter the complexity of synthetically generated datasets, a set of 10 complexity measures for classification and 4 for regression tasks has been optimized in the proposed epco approach. The proposed generic algorithm searches the space of feature projection matrices, which are used to transform the source data distribution towards a target complexity profile. The method was used to generate $550$ classification datasets and $250$ regression datasets towards five different target difficulty levels. Experiments evaluating state-of-the-art classifiers and regression methods showed a decrease in classification accuracy and an increase in regression errors when faced with datasets of higher complexity.
The presented algorithm used a custom genetic approach to many-criterion optimization, where individuals in the population were sorted according to fitness to the considered criteria and their sum. This simple approach yielded promising optimization results, especially given a small population and up to ten considered criteria. Future work may employ more sophisticated optimization strategies, in which the final individuals of the population form a more continuous Pareto frontier of the optimized complexity measures. Studying more complex transformation strategies is also a promising future research direction.